I need more help on this problem.
I got the answer from my professor and he said it was 1/8. I don't know how to get the fraction as the answer. Any help would be great!
The scruffy dog company sold 5600 dog booties last year and 4800 dog booties this year. By what FRACTION did their sales decrease this year as compared to last year?
Thanks!

Answer by stanbon(75887)

The scruffy dog company sold 5600 dog booties last year and 4800 dog booties this year. By what FRACTION did their sales decrease this year as compared to last year?
--------------------
Quantity Decrease: 5600-4800 = 800
Fractional Decrease: 800/5600 = 1/7
---------------------------
Cheers,
Stan H.